# vAttention: Dynamic Memory Management for Serving LLMs without PagedAttention

### v2

@@ 2.2. Fragmentation and PagedAttention
3. Issues with the PagedAttention Model
Despite being inspired by demand paging, the PagedAttention approach is different from it: PagedAttention mandates an implementation of paging in user space whereas conventional demand paging is transparent to applications. This section elaborates on issues that arise with such an approach.
3.1. Requires Re-writing the Attention Kernel
PagedAttention necessitates re-writing the attention kernel. This is because conventional implementations of the attention operator assume that the two input tensors K and V (Equation 2) are stored in contiguous memory. By departing from the conventional memory layout, PagedAttention requires an implementation of the attention operator to be modified so as to compute attention scores over non-contiguous KV-cache blocks. Writing correct and performant GPU kernels can be challenging for most programmers (fa-paged-crash, ).
Being a fundamental building block of the transformer architecture, the attention operator has witnessed a tremendous pace of innovation in the systems and ML communities for performance optimizations (flashattention, ; flashattention2, ; flashdecoding, ; flashdecoding++, ; flashinfer, ; h2o2023, ; flashattnhopper, ; groupedqueryattention, ; multiqueryattention, ; generatinglongsequence2019, ; longformer, ), and this trend is likely to continue. In the PagedAttention model, keeping up with new research requires continued efforts in porting new optimizations to a PagedAttention-aware implementation. Production systems can therefore easily fall behind research, potentially losing performance and competitive advantage. To provide an example, Table 6 shows that the paged kernel of vLLM is already up to 2.8×2.8$\times$2.8 × slower than the FlashAttention kernel (groupedqueryattention, ).
@@ 3.2. Adds Redundancy in the Serving Framework
PagedAttention makes an LLM serving system responsible for managing the mappings between KV-cache and dynamically allocated memory blocks. For example, consider a request that allocates four KV-cache blocks over time (left half of Figure 1). These blocks are usually non-contiguous in virtual memory. During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV-cache blocks. To facilitate this, the serving system needs to track the virtual memory addresses of KV-cache blocks and pass them to the attention kernel at runtime. This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half in Figure 1).
Figure 2. Overhead of PagedAttention in prefill kernels (model: Llama-3-8B, one A100 GPU). Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention (FA) and FlashInfer (FI).
@@ 3.3.1. Runtime overhead on the GPU
PagedAttention slows down attention computation by adding extra code in the critical path. For example, vLLM acknowledges that their PagedAttention-based implementation was $20-26\%$ slower than the original FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (vllmsosp, )).
In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries. For example, PagedAttention based prefill kernels of FlashAttention and FlashInfer are up to 28% and 24% slower than the non-paged kernels in the corresponding libraries. Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels.
To highlight another example of difficulty involved in writing an efficient attention kernel with paging support, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with larger block sizes of 64 and 128. Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in L1 cache.
@@ 3.3.2. Runtime overhead on the CPU
To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel. In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV-cache blocks in the longest request of the batch. This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV-cache blocks in each request by padding unoccupied slots with zeros. If a batch contains a few long and many short requests, such padding results in a significant overhead. In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations. While a recent fix (vllm-blocktable-issue, ) has mitigated some of this overhead, we find that it can still be as high as $10\%$. High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11%, from 412 tokens/sec to 365 tokens/sec (tensorrt-perf-decay, ). This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead. However, doing so would be an unwelcome change for most programmers.
Overall, this section shows that the PagedAttention model adds a significant programming burden while also being inefficient. vAttention introduces a more systematic approach to dynamic KV-cache memory management by leveraging the existing system support for demand paging. However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads from a memory management perspective.
@@ 4. Insights into LLM Serving Systems
To highlight the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism. We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §5.4 for our discussion and optimizations for the prefill phase).
Observation-1: On a per-iteration basis, KV-cache memory requirement is known in advance. This is a dream property for a memory allocator, thanks to the auto-regressive decoding wherein every forward pass of the model generates one token per request. Therefore, with every iteration, the KV-cache memory footprint of a request grows uniformly by one token as long as the request is active.
Observation-2: KV-cache does not require high memory allocation bandwidth. The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory. For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively. Further, each iteration runs for 10s-100s of milliseconds implying that a request requires at most a few megabytes of memory per second. While batching improves system throughput (sarathi2023, ; sarathiserve2024, ; patel2023splitwise, ; distserve2024, ), the number of tokens generated per second plateaus beyond a certain batch size (Figure 4, left). This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 128 for Yi-34B). For all the models we studied, we observe that the highest memory allocation rate is at most 600MB per second (Figure 4, right).
Figure 4. Decode throughput (left) and the rate of physical memory allocation (right) saturate at large batch sizes.
In vAttention, we leverage these observations to implement an efficient dynamic memory management system for KV-cache. In the next section, we begin with a high-level overview of vAttention (§5.1), then discuss how vAttention is used for serving LLMs (§5.3) and finally describe our optimizations (§5.4).
@@
To facilitate dynamic allocation of physical memory to unmodified attention kernels, vAttention leverages system support for demand paging instead of implementing it in user space.
@@ 5.1. Design Overview
vAttention builds on the ability to allocate virtual memory and physical memory separately. Specifically, we allocate a large contiguous buffer for the KV-cache in virtual memory ahead-of-time (similar to reservation-based allocators) while deferring the allocation of physical memory to runtime (similar to PagedAttention). This way, vAttention preserves virtual contiguity of KV-cache without wasting physical memory. This approach is feasible because memory capacity and fragmentation are limiting factors only for physical memory whereas virtual memory is abundant e.g., modern 64-bits systems provide a 128TB user-managed virtual memory per process11164-bits systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space..
Figure 5. Dynamic memory management in vAttention for a single K-cache (or V-cache) tensor. (a) shows a virtual tensor for a batch of two requests with no physical memory allocation yet. (b) R1 is allocated one physical page. (c) R1 is allocated two pages and R2 is allocated one page. (d) R1 has completed but vAttention does not reclaim its memory (deferred reclamation). (e) when R3 arrives, vAttention assigns R1’s tensor to it which is already backed by physical memory.
Pre-reserving virtual memory
Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV-cache of the maximum batch size (configurable) that needs to be supported. In doing so, we assume that each request would generate as many tokens as the maximum sequence length supported by the model.
Number of virtual memory buffers
A serving framework typically maintains separate K and V tensors for each layer of the model: we refer to them individually as K-cache and V-cache. We allocate separate virtual memory buffers for K-cache and V-cache. For a single GPU job, this requires pre-reserving $2\times N$ buffers where $N$ is the number of layers in the model. In a multi-GPU job, each worker reserves $2\times N^{\prime}$ buffers where $N^{\prime}$ is the number of layers managed by that worker ($N^{\prime}=N$ with tensor-parallelism whereas $N^{\prime}<N$ with pipeline-parallelism).
@@
Table 2. CUDA VM (virtual memory) APIs. * represents APIs that we use once while instantiating or terminating the serving framework. Rest of the APIs are used for (un)mapping physical memory pages at runtime. CUDA APIs (prefixed with cu) support only 2MB allocation sizes, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
Size of a virtual memory buffer
The maximum size of a buffer is $BS=B\times L\times S$ where B is the maximum batch size, L is the maximum context length supported by the model, and $S$ is the size of a single token’s per-layer K-cache (or V-cache) on a worker. Further, $S=H\times D\times P$, where $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). Note that $S$ is constant for a given model configuration.
Consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2). In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$. For this model, the maximum size of K-cache (or V-cache) per-worker per-layer is $S=200MB$ $(200K*4*128*2)$. Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$). Therefore, the total virtual memory requirement for 60 layers is 120 buffers of 100GB each (12TB total). Note that the amount of virtual address space available grows with the number of GPUs e.g., with two TP workers, the amount of virtual address space available is 256TB. Therefore, virtual memory allocations can be satisfied easily. Figure 5 shows how vAttention allocates physical memory pages dynamically.
5.2. Leveraging Low-level CUDA Support
The standard GPU memory allocation interface cudaMalloc does not support demand paging i.e., it allocates virtual memory and physical memory at the same time. However, recent CUDA versions provide programmers a fine-grained control over virtual and physical memory (cudavirtualmemory, ; gmlakeasplos24, ). We leverage these low-level APIs in vAttention.
@@ 5.2.1. CUDA virtual memory APIs
Table 2 provides an overview of CUDA APIs that allow separating the allocation of virtual memory from physical memory. The allocation granularity depends on the page size used by the GPU. Further, the size of a virtual memory buffer or a physical memory handle must be a multiple of the physical memory allocation granularity. It is also important to highlight that physical memory pages can be allocated (or de-allocated) to sub-regions in a virtual memory buffer independently of other sub-regions. For simplicity, we refer to the granularity of physical memory allocation as page size.
@@ 5.2.2. Extending PyTorch caching allocator
KV-cache is a collection of tensors. In current deep learning frameworks such as PyTorch, a tensor allocated via APIs such as torch.empty comes with pre-allocated physical memory. This is because the PyTorch caching allocator uses the cudaMalloc interface to allocate GPU memory (both virtual and physical). Relying on the low-level API support from CUDA, we extend the PyTorch caching allocator to allow an application to reserve a virtual memory buffer for a tensor without committing physical memory ahead-of-time. We refer to tensors allocated via these APIs as virtual tensors.
5.2.3. Request-level KV-cache indexing
Note that each virtual tensor represents the K-cache (or V-cache) of a layer for the maximum batch size B. In these tensors, different requests occupy different non-overlapping sub-regions (say sub-tensors). We locate the sub-tensor of a request with a unique integer identifier reqId that lies in the range of 00 to $B-1$ (note that at most $B$ requests run simultaneously). The K-cache (or V-cache) offset of a request’s sub-tensor in the virtual tensor of the entire batch is $\texttt{reqId}\times S$ where $S$ is the maximum K-cache (or V-cache) size of a request on a worker. The request identifier reqId is allocated by vAttention.
5.3. Serving LLMs with vAttention
We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers. Our library exposes a set of simple APIs to the serving framework (see Table 3 and Algorithm 1).
5.3.1. Initial setup
When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N^{\prime},H,D,P$, $B$ and a preferred page size via the init API. Internally, vAttention reserves $2\times N^{\prime}$ virtual tensors (using our modified PyTorch caching allocator) for the KV-cache at each worker. These virtual tensors are reserved for the lifetime of the serving application. In addition, vAttention also pre-allocates physical memory pages during initialization. However, these pages are not mapped into the KV-cache yet.
5.3.2. Scheduling a new request
When a new request is scheduled for the first time, the serving framework obtains a new reqId from vAttention via alloc_reqid. All subsequent memory management operations of the request are tagged with this reqId.
@@ 5.3.2. Scheduling a new request
arguments: $N^{\prime},B,L,H^{\prime},P$, page_size.
return value: a list of KV-cache tensors.
@@
step Ensures physical memory pages are mapped
@@
Table 3. Key APIs that vAttention exposes to a serving framework for dynamic KV-cache memory management.
5.3.3. Model execution
Before scheduling a batch for execution, the framework needs to ensure that the KV-cache sub-tensors of each active request are backed by physical memory. For this purpose, before dispatching the first kernel of an iteration to the GPU, the framework invokes the step API, specifying the current context length of each request (context length is set to 0 for each inactive reqId). Internally, vAttention ensures that enough physical pages are mapped for each active reqId before returning execution back to the framework. If vAttention cannot satisfy the memory demand, it returns with a failure in response to which a serving framework can preempt one or more requests to allow forward progress (this is similar to vLLM’s default behavior). We leave more sophisticated policies such as swapping out KV-cache to CPU memory as future work.
Depending on whether a request is in the prefill phase or decode phase, different number of physical memory pages may need to be mapped for a given iteration. The prefill phase processes the input tokens of given prompt in parallel and populates one slot in the K-cache (and V-cache) of the request at each layer of the model. Therefore, the number of pages needed to be mapped depends on the number of prompt tokens being scheduled. If the total K-cache size of all prompt tokens at one layer of the model is $s$ and page size is $t$, then each worker needs to ensure that at least $(s+t-1)/t$ physical memory pages are mapped in each of the $2\times N^{\prime}$ KV-cache sub-tensors of the given reqId.
For a request in the decode phase, the number of new pages required is at most one per request. This is because each iteration produces only one output token for a request. vAttention internally tracks the number of pages mapped for each request and maps a new page only when the last page allocated to that request is fully utilized.
5.3.4. Request completion
A request terminates when a user specified context length or the maximum context length supported by the model is reached, or when the model produces a special end-of-sequence token. The framework notifies vAttention of a request’s completion with free_reqid. Internally, vAttention may unmap the pages of a completed request or defer them to be freed later.

### v3

@@ 3.2. Adds Redundancy in the Serving Framework
PagedAttention makes an LLM serving system responsible for managing the mappings between KV cache and dynamically allocated memory blocks. For example, consider a request that allocates four KV cache blocks over time (left half of Figure 1). These blocks are usually non-contiguous in virtual memory. During the computation of Equation 2, PagedAttention kernel needs to access all the elements of the four KV cache blocks. To facilitate this, the serving system needs to track the virtual memory addresses of KV cache blocks and pass them to the attention kernel at runtime. This approach effectively requires duplicating what the operating system already does for enabling virtual-to-physical address translation (right half in Figure 1).
@@ 3.3.1. Runtime overhead on the GPU
PagedAttention slows down attention computation by adding extra code in the critical path of execution. For example, the vLLM paper acknowledges that the PagedAttention-based implementation was $20-26\%$ slower than the corresponding none-paged FasterTransformer kernel, primarily due to the overhead of looking up Block-Tables and executing extra branches (see Figure 18a in  (Kwon et al., 2023)). In addition, Figure 2 shows that incorporating PagedAttention has also added a significant performance overhead in other state-of-the-art kernel libraries. For example, PagedAttention based prefill kernels of FlashAttention-2 and FlashInfer are up to 37% and 42% slower than the non-paged kernels in the corresponding libraries. Our analysis reveals that the number of instructions executed in PagedAttention kernels is $7-13\%$ higher than the non-paged kernels. Caching page indices also increases register pressure, causing register spilling.
To highlight another example of difficulty involved in writing an efficient paged kernel, Figure 3 shows that the performance of vLLM’s paged decode kernel is significantly worse with large block sizes of 64 and 128. Our analysis indicates that this is likely due to L1 cache efficiency: smaller blocks have a higher memory bandwidth utilization due to higher hit rates in the L1 cache.
Figure 2. Overhead of PagedAttention in prefill kernels (model: Llama-3-8B, one A100 GPU). Numbers on top show overhead over the corresponding non-paged implementation of FlashAttention-2 (FA2) and FlashInfer (FI).
@@ 3.3.2. Runtime overhead on the CPU
To enable PagedAttention, a serving system needs to supply Block-Tables to the attention kernel. In vLLM, the latency of preparing a Block-Table depends on batch composition and grows proportional to max_num_blocks $\times$ batch_size where max_num_blocks refers to the number of KV cache blocks in the longest request of the batch. This is because vLLM manages a Block-Table as a 2D tensor and aligns the number of KV cache blocks in each request by padding unoccupied slots with zeros. If a batch contains a few long and many short requests, such padding results in a significant overhead. In our earlier experiments, we observed that Block-Table preparation in vLLM was contributing 30% latency in decode iterations. While a recent fix (vll, 2024b) has mitigated some of this overhead, we find that it can still be as high as $10\%$. High overhead of PagedAttention has also been found in TensorRT-LLM, degrading throughput by 11% (ten, 2023). This issue was attributed to the Python runtime of TensorRT-LLM and moving to a C++ runtime can mitigate the CPU overhead. However, doing so requires non-trivial programming effort.
Overall, this section shows that PagedAttention adds a significant programming burden while also being inefficient. In vAttention, we propose a more principled approach to dynamic KV cache memory management. However, before delving into vAttention, we first highlight some of the fundamental characteristics of LLM serving workloads from a memory management perspective.
@@ 4. Insights into LLM Serving Systems
To understand the memory allocation pattern of LLM serving systems, we experiment with Yi-6B running on a single A100 GPU, and Llama-3-8B and Yi-34B running on two A100 GPUs with tensor-parallelism (TP). We set the initial context length of each request to 1K tokens, vary the batch size from 1 to 320 and measure the throughput and memory requirement of the decode phase (see §6 for our discussion and optimizations for the prefill phase).
Observation-1: KV cache memory requirement is predictable on a per-iteration basis. Due to auto-regressive decoding, once a request enters the decode phase, its KV cache size increases uniformly by one token per iteration. This allows the serving system to determine in advance if additional memory will be required during the iteration’s execution. [^1] Note that this prediction is possible only at the granularity of individual iterations; the total KV cache requirement of a request remains unknown as it depends on the total number of output tokens in the request.
Observation-2: KV cache does not require high memory allocation bandwidth. The memory footprint of a single token across all layers is typically few 10s-100s of kilobytes of memory. For example, the per-token memory footprint of Yi-6B, Llama-3-8B and Yi-34B is 64KB, 128KB and 240KB, respectively. Further, each iteration runs for 10s-100s of milliseconds implying that a request requires at most a few megabytes of memory per second. While batching improves system throughput (Agrawal et al., 2023, 2024b; Patel et al., 2024; Zhong et al., 2024), the number of tokens generated per second plateaus beyond a certain batch size (4(a)). This implies that the memory allocation bandwidth requirement also saturates at large batch sizes (e.g., at 256 for Yi-34B). For all the models we studied, we observe that the highest memory allocation rate is at most 750MB per second (4(b)). vAttention leverages these observations to optimize KV cache memory management.
(a) Decode throughput.
(b) Rate of memory allocation.
Figure 4. Decode throughput (top) and the rate of physical memory allocation (bottom) saturate at large batch sizes.
Figure 5. Dynamic memory management in vAttention for a single K cache (or V cache) tensor. (a) shows a virtual tensor for a batch of two requests with no physical memory allocation yet. (b) R1 is allocated one physical page. (c) R1 is allocated two pages and R2 is allocated one page. (d) R1 has completed but vAttention does not reclaim its memory (deferred reclamation). (e) when R3 arrives, vAttention assigns R1’s tensor to it which is already backed by physical memory.
@@
Our primary observation is that physical memory fragmentation can be avoided without making KV cache non-contiguous in virtual memory. To realize this, vAttention decouples the allocation of virtual memory from physical memory by leveraging system support for demand paging (instead of implementing demand paging in user space, as in PagedAttention).
@@ 5.1. Design Overview
vAttention employs distinct allocation policies for virtual and physical memory. Specifically, we allocate a large contiguous buffer for the KV cache in virtual memory ahead-of-time (similar to systems prior to PagedAttention) while deferring the allocation of physical memory to runtime (similar to PagedAttention). This design preserves virtual contiguity of KV cache without fragmenting physical memory. Note that this approach could fragment and waste virtual memory. However, this is not an issue since virtual memory is abundant, e.g., modern 64-bit systems provide a 128TB user-addressable virtual memory per process.22264-bit systems use only 48 bits for virtual addresses today, providing a per-process virtual memory space of 256TB which is divided equally between the user space and (OS) kernel space.
5.1.1. Pre-reserving virtual memory
Since virtual memory is abundant, we pre-allocate it in size that is large enough to hold the KV cache of the maximum batch size (configurable) that needs to be supported. In doing so, we assume that each request’s context length is same as the maximum supported by the model.
5.1.2. Number of virtual memory buffers
A serving framework maintains separate K and V tensors for each layer of the model. Therefore, we reserve $2\times N$ buffers on a worker where $N$ is the number of layers managed by that worker.
@@ 5.1.2. Number of virtual memory buffers
Table 3. CUDA VMM APIs. * represents APIs that we use once while instantiating or terminating the serving framework. Rest of the APIs are used for (un)mapping physical memory pages at runtime. CUDA APIs (prefixed with cu) support only 2MB pages, whereas our CUDA extension APIs (prefixed with v) support fine-grained allocations.
5.1.3. Size of a virtual memory buffer
The maximum size of a buffer is $BS=B\times S$ where B is the maximum batch size and $S$ is the maximum size of a single request’s per-layer K cache (or V cache) on a worker. Further, $S=L\times H\times D\times P$, where $L$ is the maximum context length supported by the model, $H$ is the number of KV heads on a worker, $D$ is the dimension of each KV head and $P$ is the number of bytes based on model precision (e.g., P=2 for FP16/BF16). As an example, consider Yi-34B with FP16 and two-way tensor-parallelism (TP-2). In this case, $N\!=\!60,H=4,D=128,P=2$ (8 KV heads of Yi-34B are split evenly on two GPUs), and maximum supported context length $L=200K$. For this configuration, $S=200MB$ $(200K*4*128*2)$. Assuming $B=500$, the maximum size of each buffer per-worker is $BS=100GB$ ($500\times 200MB$). Therefore, the total virtual memory requirement for 60 layers is 120 buffers of 100GB each (12TB total). Note that the size of virtual address space available grows with the number of workers, e.g., with two TP workers, the total size of user-addressable virtual address space is 256TB. Therefore, virtual memory is always plentiful to satisfy large allocations. Figure 5 shows an example of how vAttention allocates physical memory pages dynamically.
5.2. Leveraging CUDA Virtual Memory Support
The standard GPU memory allocation interface cudaMalloc does not support demand paging, i.e., it allocates virtual memory and physical memory at the same time. However, recent CUDA versions provide programmers a fine-grained control over managing virtual and physical memory, including support for decoupling their allocations (cud, 2024; Guo et al., 2024). We leverage these low-level APIs.
@@ 5.2.1. CUDA virtual memory APIs
Table 3 provides an overview of CUDA VMM APIs that allow decoupling the allocation of virtual memory from physical memory. The allocation granularity depends on the page size used by the GPU. Further, the size of a virtual memory buffer or a physical memory handle must be a multiple of the physical memory allocation granularity. Physical memory pages can be allocated to (or de-allocated from) sub-regions in a virtual memory buffer independently of other sub-regions.
Figure 6. An LLM serving system interacts with vAttention for KV cache management with a set of simple APIs listed in Table 4. All other memory objects (e.g., activations) are allocated by the standard PyTorch caching allocator.
@@ 5.2.2. Extending PyTorch caching allocator
KV cache is a collection of tensors. In current deep learning frameworks such as PyTorch, a tensor allocated via APIs such as torch.empty comes with pre-allocated physical memory. This is because the PyTorch caching allocator relies on the cudaMalloc interface (Figure 6). Relying on the low-level API support from CUDA, we extend the PyTorch caching allocator to allow an application to reserve a virtual memory buffer for a tensor without committing physical memory ahead-of-time. We refer to tensors allocated via these APIs as virtual tensors.
5.2.3. Request-level KV cache indexing
A virtual tensor represents the K cache (or V cache) of a layer for the maximum batch size B. In these tensors, different requests occupy different non-overlapping sub-regions (say sub-tensors). We locate the sub-tensor of a request with a unique integer identifier reqId that lies in the range of 00 to $B-1$ (note that at most $B$ requests run simultaneously). The K cache (or V cache) offset of a request’s sub-tensor in the virtual tensor of the entire batch is $\texttt{reqId}\times S$ where $S$ is the maximum size of per-layer K cache (or V cache) of a request on a worker. The request identifier reqId is allocated by vAttention.
@@ 5.2.3. Request-level KV cache indexing
arguments: $N,B,L,H,D,P$, page-group-size.
return value: a list of KV cache tensors.
@@
step Ensures KV cache tensors are backed by physical pages up to the current context length of each active request
@@
Table 4. Key APIs that vAttention exposes to a serving framework for dynamic KV cache memory management.
5.3. Serving LLMs with vAttention
We build vAttention as a Python library that internally uses a CUDA/C++ extension for interacting with CUDA drivers. Our library exposes a set of simple APIs to the serving framework (shown in Figure 6Table 4 and Algorithm 1). For simplicity, we discuss the use of these APIs from a single worker’s perspective; all workers behave the same.
5.3.1. Initial setup
When the serving framework starts, each model worker loads the vAttention library and configures it with model parameters $N,H,D,P$, $B$ and a preferred page-group size (§6.2) via the init API (line 4 in Algorithm 1). Internally, vAttention reserves $2\times N$ virtual tensors on the worker, as shown in Figure 5(a), where $N$ is the number of layers hosted by the worker. These virtual tensors are reserved for the lifetime of the serving application. In addition, vAttention also pre-allocates physical memory pages at each worker during initialization. However, these pages are not mapped into the KV cache at this point.